A ball is thrown upward with a velocity 15 m/s. What will be the ball’s velocity when it reaches a height of 0.75 m above the ground?
Answer #1

Using equation of motion,

v^2=u^2+2gs

v=final velocity

u=initial velocity=15m/s

g=gravitational acceleration=10m/s^2

s=distance travelled= 0.75m

v^2=(15)^2+2×10×0.75

v^2=225+150

v^2=375

v=√375

v=19.37m/s

Answer is approximately v=~19.4m/s